OpenStudy (anonymous):

I meant to say multiply the two binomials below to produce a difference of squares....(2A+3B)(2A-3B)

OpenStudy (bahrom7893):

4A^2 - 9B^2

OpenStudy (saifoo.khan):

4A^2 - 9B^2

OpenStudy (anonymous):

Yes ^

OpenStudy (anonymous):

Yeah, I posted that in the other question

OpenStudy (anonymous):

identity (a - b) * (a + b) = a² - b² So, (2a+3b)(2a-3b) = (2a)² - (3b)² = 4a² - 9b² => 4a to the second power - 9b to the second power